OpenStudy (anonymous):

Evaluate the expression 1/2 a^2 b for a=2/7 b=1/2

OpenStudy (anonymous):

@DavidUsa

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

=(1/2)(2/7)=1/7 because a^2b = a WHEN b=1/2